Standard supper market issued $8,000,000 of 5-year, 8% callable bonds dated July 1, 2013, at an effective rate of 10%. Interest is payable semiannually on December 31 and June 30.

Required:

Calculate the present amount of the bonds.

Prepare a journal entry to record the issuance of bonds.

Determine the total interest expense for the bonds

Prepare an amortization table using the straight-line method.


1
Expert's answer
2021-07-05T08:57:28-0400

Solution:

a.). The present value of the bonds = The present value of a bond’s interest payments + the present value of a bond’s maturity amount.

First, calculate the present value of a bond’s interest payments:

Present Value of an Ordinary Annuity (PVOA) = PMT x PVOA factor

Bond interest rate = 8%

Effective interest rate = 10% = 10%/2 = 5% semi-annually

Interest payments to be paid over the 5-year period = 8,000,000 x 8% x 6/12 = 320,000

To calculate the present value of the semi-annual interest payments of $320,000 each, you need to discount the interest payments by the effective interest rate for the six-month period.

 

PVOA factor ="[\frac{1 - (1 + r)^{-n}}{r }] = [\frac{1 - (1 + 0.05)^{-10}}{0.05 }] = 7.722"

Present Value of an Ordinary Annuity (PVOA) = 320,000 x 7.722 = 2,471,040

 

Calculate the present value of the bond’s maturity amount:

Present value of the bond’s maturity amount = FV x PV of 1 factor

FV = 8,000,000


PV of 1 factor = "\frac{1}{(1 + r)^{n} } = \frac{1}{(1 + 0.05)^{10} } = 0.614"

Present value of the bond’s maturity amount = 8,000,000"\times" 0.614 = 4,912,000

The present value of the bonds = 2,471,040 + 4,912,000 = 7,383,040

 

b.). Journal entry for the issuance of bonds:

Determine the discount on bonds payable = 8,000,000 – 7,383,040 = 616,960


The journal entry to record an $8,000,000 bond that was issued for $7,223,040 on July 1, 2013, is:

Dr. Cash                                                         7,383,040

Dr. Discount on bonds payable                     616,960

              Cr. Bonds payable                                                     8,000,000

              

 

c.). The total interest expense for the bonds:

The total interest expense for the bonds = "(8\%\times8,000,000 \times \frac{6}{12} \times10) = \$3,200,000" 


d.). Amortization table using the straight-line method:

Find the below amortization table: